Your answer is incorrect. Try again. Nate Diaz, a UFC fighter, has landed 1044 significant strikes in his previous 12 UFC fights. Assume that the number of significant strikes is normally distributed with ơ-10. Find the 95% two-sided confidence interval on the mean number of strikes. Round your answers to two decimal places (e.g. 98.76) 1038.34 1049.66
Answer #1

Mean number of stries=1044/12=87

\sigma=10

n=12

95% Confidence interval on the mean number of strikes is

\bar{x}\pm Z_{\alpha/2}*\frac{\sigma}{\sqrt{n}}

87\pm 1.96*\frac{10}{\sqrt{12}}

(81.34,92.66)

81.34 \leq \mu \leq 92.66